### v1

# Label-Efficient Grasp Joint Prediction with Point–JEPA

Abstract
We study whether 3D self-supervised pretraining with a Joint-Embedding Predictive Architecture (Point–JEPA) enables label-efficient grasp joint-angle prediction. Meshes are converted to point clouds and tokenized; a ShapeNet-pretrained Point–JEPA encoder feeds a lightweight multi-hypothesis head trained with winner-takes-all and evaluated by top–logit selection. On DLR–Hand II with strict object-level splits, Point–JEPA improves top–logit RMSE and Coverage@15° in low-label regimes (e.g., 26% lower RMSE at 25% data) and reaches parity at full supervision. This suggests JEPA-style pretraining is a practical lever for data-efficient grasp learning.
I Related Work
Self-supervised learning (SSL) for 3D data has largely progressed along three directions.
(i) Reconstruction-based methods learn by masking and reconstructing inputs in the input space. On point clouds this includes point/voxel masked autoencoding; e.g., Voxel-MAE reconstructs masked voxels for sparse automotive LiDAR and improves downstream tasks with fewer labels [1, 2, 3, 4].
(ii) Teacher–student (latent-target) objectives predict contextualized latent targets rather than raw inputs. data2vec provides a modality-agnostic recipe that predicts latent representations of the full input from a masked view, while Point2Vec adapts this idea to point clouds and addresses positional-leakage pitfalls unique to 3D, reporting strong transfer on ModelNet40/ScanObjectNN [5, 6].
(iii) Joint-embedding predictive architectures (JEPA) predict target representations for spatially contiguous blocks from a single context block, learning semantic features without heavy view augmentations [7, 8]. Point–JEPA brings this design to point clouds with a simple sequencer that orders patch centers so contiguous context/target blocks can be sampled despite unordered points, avoiding input-space reconstruction or extra modalities while remaining competitive on standard 3D benchmarks [9, 8].
Despite this progress, the impact of predictive, context-aware pretraining on grasp–joint prediction remains underexplored [10, 11, 12]. In this study, we adopt a Point–JEPA backbone for object embeddings and pair it with an inference-aware, multi-hypothesis loss for joint angles, comparing against training from scratch and analyzing label-efficiency gains under fixed object-level splits.
II Technical Outline
We convert meshes to point clouds and tokenize them into local patches for transformer processing [13].
We utilize a pretrained JEPA backbone trained on ShapeNet point clouds via Point–JEPA [9]. A block sampler selects spatially contiguous target windows, and a sequencer orders patch centers to keep these windows contiguous in the token sequence, following the JEPA masking/targeting scheme [7]. The context encoder sees masked input; the target encoder sees full input to produce latent targets. A lightweight predictor maps context features to the target space; we regress predictions to stop-gradient targets on masked blocks. The target encoder is an EMA copy of the context encoder; no negatives or input reconstruction are required. After pretraining we discard the predictor and EMA target, retaining the context encoder for contextualized patch features (Fig. 2); attention pooling yields a global object embedding (Fig. 3). For clarity, Fig. 2 depicts single masked patches; training uses block masking.
For grasp prediction, the pooled object embedding is concatenated with the 7D wrist pose and passed to a lightweight head that emits joint-angle hypotheses and logits. Training uses a winner-takes-all / min-over- objective to preserve multi-modality (see Sec. III-D); at inference, we select the top–logit hypothesis [14].
III Methodology
III-A Dataset & split protocol
Data and filtering. We use the DLR–Hand II dataset (non-public; for code/protocol/weights see Footnote Label-Efficient Grasp Joint Prediction with Point–JEPA ) built atop ShapeNet object categories, comprising a range from 10 to 500 unique shapes per category. Each shape provides around 240 grasp samples, each consisting of a 7D wrist pose, a 12D joint configuration, and a grasp quality score. After discarding all grasps with score , our dataset contains on the order of 1,500 meshes and 370,000 total grasps.
Object-level, category-stratified splits. To prevent overfitting and pose leakage from multiple grasps of the same mesh, we split at the object level. Sample-level splits proved overly easy and invalidated the value of SSL-based encoding, and so were not used.
Fixed evaluation. We fix one suite consistent for all label budgets: val (10% of objects) and test (another 10%).
Low–data training subsets. For , we sample the train set from the remaining objects with proportional category stratification (at least one object per synset).
Two split packs and reproducibility. We construct two independent train split packs (A/B) with different manifest seeds; validation fixtures (val, test) are identical across packs. Unless otherwise stated, entries report the mean SD over initialization seeds (0,1) and, where available, over split packs. Due to compute constraints, Pack B was run at anchor budgets (25% and 100%); the low-label budgets (1%, 10%) use Pack A. At the anchor points, Pack A and Pack B yield qualitatively similar trends, so we adopt Pack A as the default and use Pack B to check pack sensitivity.
III-B Preprocessing & object representation
Meshes are converted to point clouds and tokenized into local groups to create patch embeddings. We use points per object and a tokenizer with groups (group size , radius ). These patch embeddings feed the self–supervised encoder (Fig. 2); a global object embedding is then produced by attention pooling, which we found to be more stable and accurate than mean/max pooling.
III-C SSL backbone (Point–JEPA)
We use the Point–JEPA context encoder as the backbone (predictor/target discarded after pretraining). Tokenized point-cloud patches with positional encodings are encoded and attention-pooled into a single object embedding. We fine-tune end-to-end with a two-tier learning rate (smaller on the pretrained backbone, larger on the new grasp head); concrete settings and schedules appear in Sec. IV.
III-D Joint estimation head & loss
We predict candidate joint configurations for each object embedding and wrist pose, together with a logit vector used to rank hypotheses at test time. Grasping is inherently multi-modal: several distinct joint settings can succeed for the same pose, and single-output MSE collapses these modes to their mean, often yielding an interpolated (and infeasible) grasp (see Fig. 4). To preserve distinct modes, we use a winner-takes-all (WTA) / min-over- objective, a standard approach for ambiguous prediction tasks.
The first term penalizes only the closest hypothesis, avoiding mean-collapse; the cross-entropy term teaches the logits to act as a proxy selector for the winning mode. We then use the same selector at inference (pick the top–logit hypothesis), so evaluation does not rely on an oracle “best-of-”. This mirrors multi-hypothesis practice in trajectory forecasting, where sets of candidates are scored and one is chosen at test time.
In preliminary trials we also considered Mixture Density Networks (MDNs) to model multi-modality probabilistically, but found them brittle in this high-dimensional joint space (training instability and covariance issues) and less aligned with our top-1 selection at inference; the WTA formulation remained simpler and more stable.
All experiments use to balance coverage and compute (brief pilots with were inconclusive; a full sweep is left to future work). Beyond the primary top–logit RMSE, we report two diagnostics that probe the multi-hypothesis mechanism without relying on an oracle at evaluation time: (i) a selection gap (top–logit RMSE minus best-of- RMSE) to assess selector fidelity, and (ii) Coverage@, the fraction of samples for which at least one hypothesis lies within of the ground truth.
III-E Evaluation protocol & metrics
At inference, we choose the head with the largest logit (“top–logit”). The main metric is top–logit RMSE on joint angles. We additionally compute coverage@ (a grasp counts as covered if any head is within the angular threshold) and a diversity proxy (mean pairwise dispersion across heads) to quantify multimodality.
IV Results
All metrics are reported on the fixed test split with object-level stratification, averaged over split packs and seeds where available.
IV-A Main outcomes
Table I summarizes top–logit RMSE (radians; lower is better). Three patterns emerge:
Low labels (1–10%): JEPA pretraining yields consistent gains (– relative error reduction), indicating improved label efficiency in scarce-data regimes.
Moderate labels (25%): The largest improvement appears at 25% ( relative), a “sweet spot” where pretrained context helps most before fully supervised training saturates.
Full labels (100%): Performance is effectively at parity; differences are within seed variance (JEPA scratch), suggesting that with enough supervision, training from random initialization can catch up.
IV-B Inference-aware selection and coverage
To assess whether the logits reliably choose the correct hypothesis, we define the selection gap
with smaller values indicating better alignment between the learned selector and the oracle winner. At 1% and 10% budgets, JEPA reduces the gap relative to scratch (e.g., 1%: vs rad; 10%: vs rad), demonstrating stronger inference-aware selection. The full curves and plotting scripts are provided in our Git repository.
| Train split | Scratch | JEPA | (rel.) |
| 1% | 0.363 0.002 | 0.335 0.003 | +7.7% |
| 10% | 0.335 0.003 | 0.303 0.009 | +9.6% |
| 25%† | 0.332 0.002 | 0.246 0.012 | +25.9% |
| 100%† | 0.235 0.002 | 0.234 0.008 | +0.4% |
As a complementary multi-hypothesis metric, we report Coverage@: a grasp is covered if any head lies within of the ground-truth joint angles. Coverage increases with label budget and is consistently higher with JEPA in the low-label settings (e.g., at 10%: vs ; at 1%: vs ); see the repository for the corresponding curves and scripts.
IV-C Sensitivity to learning rates (10% grid)
We ran a 32 grid over on the 10% split (time constraints precluded repeating per budget). Top–logit RMSE versus training step shows a broad plateau at convergence: final differences across the grid are small ( rad). Larger backbone LRs destabilize early fine-tuning, while smaller head LRs slow convergence. We therefore fix for all budgets.
V Conclusion
We investigated whether 3D self-supervised pretraining improves grasp-joint prediction under limited labels. Integrating a Point–JEPA backbone with a winner-takes-all multi-hypothesis head yields consistent gains in low-label regimes on DLR–Hand II, with the largest relative improvement at 25% data. At 100% labels, training from scratch attains parity. Diagnostics indicate that the selector is effective (smaller selection gap) and that Coverage@ improves with pretraining, supporting the inference-aware design.
Results indicate JEPA-style pretraining is a practical label-efficiency lever for grasping systems when annotation is expensive.
VI Future work
We outline several directions to strengthen generalization and deployment relevance:
Cross-domain evaluation. Run the pre-registered test_category suite and real-robot tests to assess transfer beyond object-level splits.
Heads and selection. Sweep and study diversity-encouraging variants, while monitoring selection gap and top–logit coverage as primary inference-aware metrics.
Geometry-aware patching. Replace FPS + fixed-kNN with patches built from simple geometric cues, e.g., multi-scale radius neighborhoods, curvature-guided grouping, and small overlaps, so each patch better respects local shape.
Geometry-focused tokenizer. Swap the PointNet patch encoder for one that uses explicit local geometry. The JEPA objective stays the same; only the patch encoder changes.
Compute-efficient fine-tuning. Evaluate lightweight adaptation (e.g., LoRA) and backbone LR schedules across label budgets.